# First Study of the Nuclear Response to Fast Hadrons via Angular Correlations between Pions and Slow Protons in Electron–Nucleus Scattering

### v2

@@ Abstract
We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus (eAeA) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark. The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets. For heavier nuclei, the pion-proton correlation function is more spread-out in azimuth than for lighter ones, and this effect is more pronounced in the $\pi p$ channel than in earlier $\pi\pi$ studies. The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets. These trends are qualitatively reproduced by state‑of‑the‑art eAeA event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing. At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in eAeA scattering.
@@ 1 Introduction
The production and propagation of hadrons through nuclei remain central topics in quantum chromodynamics (QCD). Key questions include: “How does the process of hadronization manifest within a nucleus?”, and “How are different hadrons produced in a single scattering event correlated with each other?” LRP2023. Addressing these questions calls for detailed studies of diverse reaction channels, encompassing multiple hadron flavor combinations in the final state and spanning a broad kinematic range, to probe the different timescales of hadronization.
Despite extensive theoretical and phenomenological work on nucleon production in nuclear deep-inelastic scattering (DIS) E665:1994aiu, Degtyarenko:1997, BEBCWA59:1989ayi, CiofidegliAtti:2004pv, Palli:2009it, Larionov:2018igy, PhysRevD.100.073010, Chang:2022hkt, RoblesGajardo:2022efe, no dedicated study has correlated the production of low-energy protons with the kinematics of leading hadrons, whose spectra are known to be modified by cold nuclear matter effects as observed by the HERMES HERMES:2000ytc, HERMES:2003icw, Airapetian:2007vu, Airapetian:2009jy, Airapetian:2011jp and CLAS CLAS:2021jhm, CLAS:2022oux, CLAS:2012tlh, kbhz-h4jv collaborations. The present work addresses this gap by employing pion-proton correlations over a wide azimuthal and rapidity range, thereby linking studies of current and nuclear fragmentation in DIS off deuterium and heavy nuclei.
In this study, we examine electron–nucleus (eAeA) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig. 1. Correlating these two hadrons provides new insight into how the nucleus responds to a fast quark or hadron and how the hadronization process is modified within the nucleus. Most of the observed protons come from fragmentation of the nuclear remnant. Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum. This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements PhysRevC.111.035201.
@@
Protons produced through nuclear fragmentation, historically referred to as “grey” protons, were investigated in earlier scattering experiments with both hadronic and leptonic beams E910:1999ozb, E665:1994aiu, Degtyarenko:1997. Measurements of slow protons, together with model calculations, served to estimate the collision centrality, expressed as the mean number of projectile–nucleon interactions. Lepton measurements offer a cleaner test of intra-nuclear hadron-cascade models; in hadron collisions, multiple projectile–nucleon interactions complicate particle‐production mechanisms. Similar studies were carried out with neutrino DIS on nuclear targets BEBCWA59:1989ayi, although with much more limited precision.
Ciofi degli Atti et al. CiofidegliAtti:2004pv, Palli:2009it proposed that the dependence of slow-proton production over a wide kinematic range can serve to discriminate among hadronization models, and their calculation reproduced well the data from Ref. E665:1994aiu. Larionov et al. Larionov:2018igy showed that slow-neutrons provide similar insight into intranuclear cascades.
The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in eAeA collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
@@ 2 Experimental setup
The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters ($C$), and an electromagnetic calorimeter (EC). Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter. The drift chambers functioned in a toroidal magnetic field up to 2 Tm, where the magnet’s polarity made negatively charged particles bend toward the beam pipe. The resulting resolutions for charged particles were $\sigma_{p}/p=0.5\%$ for the momentum, 2 mrad for the polar angle, and 4 mrad for the azimuthal angle. The data were selected with a trigger that required at least one electron candidate with momentum $p>500$ MeV/c/c.
@@ 3 Event selection and observables
Following Refs. CLAS:2012tlh, CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC. Background from $\pi^{-}$ was suppressed to the $1\%$ level using the $C$ and the EC. Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs. The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the $C$ to suppress proton background. Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution. The protons’ energies were corrected for energy loss in the target using the same procedure as was used in Ref. CLAS:2022oux.
Similar to Refs. CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, we selected events with $Q^{2}>1$ $2W>2 GeV/c2c^{2}$ and $2.3<\nu<4.2$ GeV. Here, $Q^{2}$ is the square of the four-momentum transfer from the electron to the struck nucleon,
@@
is the magnitude of the four-momentum of the virtual photon+struck nucleon system, where $m_{p}$ is the mass of a proton, $\nu=E-E^{\prime}$ is the energy transfer, and $E$ and $E^{\prime}$ are the beam- and scattered-electron energies.
We selected events with a “leading” pion, defined as having fractional energy $z_{1}=E_{h}/\nu>0.5$ HERMES:2005mar, CLAS:2022asf, PhysRevC.111.035201, where $E_{h}$ is the energy of the pion. Following Refs. CLAS:2022oux, Hen:2014nza, we required the proton to have momentum greater than 350 MeV/c/c so that it could be well-reconstructed. The maximum proton momentum in our accepted $\pi p$ event sample was about 2.7 GeV/c/c. In order to ensure that the azimuthal angles of both particles are well defined, we further required that both the proton and the leading pion have transverse momenta (defined with respect to the momentum-transfer vector $\vec{q}$) greater than 70 MeV/c/c.
@@
where $E_{h}$ is the hadron’s energy and $p_{z,h}$ is its momentum in the momentum-transfer direction.
We selected particles arising from scattering from either the deuterium or nuclear targets by using the longitudinal vertex position defined by intersecting their trajectories with the beamline. The resulting vertex resolution ensured negligible ambiguity in the target tagging CLAS:2021jhm.
@@
where $N_{e^{\prime}\pi}$ is the number of events with a scattered electron and a leading $\pi^{+}$, and $\Delta\phi\equiv\phi_{\pi}-\phi_{p}$ is the azimuthal separation between the pion and proton, as defined around the momentum-transfer axis. We determined the expression in parentheses by taking $N_{e^{\prime}\pi p}$, the number of events with a scattered electron, a leading $\pi^{+}$ and a proton in each bin in $|\Delta\phi|$ and $\Delta Y$, and dividing that by the bin widths in both variables ($\pi/8$ and 0.5, respectively), and by a factor of two, since each $|\Delta\phi|$ bin covers two $\Delta\phi$ bins. These correlation functions are corrected for pair-acceptance effects using Monte Carlo (MC) simulations. Following Ref. PhysRevC.111.035201, we also used a data-driven method to correct for contamination of the deuterium sample due to scattering off the aluminum endcap walls of the liquid deuterium cells. Details for both of these corrections can be found in the Supplementary Materials supplementary.
@@
$C_{0}=\frac{N^{D}_{e^{\prime}\pi}}{N^{D,\rm tot}_{e^{\prime}\pi p}}$ (4)
where $N^{D}_{e^{\prime}\pi}$ is the number of events from the deuterium dataset with the scattered electron and leading pion, and $N^{D,\rm tot}_{e^{\prime}\pi p}$ is the total number of events in the deuterium dataset with the scattered electron, leading pion, and proton, for all kinematic bins.
From here, we determined the nuclear-to-deuterium ratio of the correlation functions, $R=C_{A}/C_{D}$, which can be used to compare the correlation functions of the nuclear targets with those of deuterium.
We determined from these correlation functions the azimuthal width $\sigma$, defined by PhysRevC.111.035201
@@
as well as the azimuthal broadening PhysRevC.111.035201 (using carbon as our reference target [^2] We chose carbon instead of deuterium (as was done in Ref. PhysRevC.111.035201) for the reference target here, since the results for the nuclear targets are more similar to one another than they are to deuterium.),
@@ 4 Uncertainties
We performed studies on various possible sources of systematic uncertainties using data and simulation studies. For the simulation studies we used the PYTHIA6 event generator and the GSIM package GSIM, which is based on GEANT3 Brun:1994aa [^3] It should be noted that Geant4 GEANT4:2002zbu had not been released yet when GSIM was being developed, and that GEANT3 was still the standard at the time that the data were taken., to simulate the response of the CLAS detector and dual-target setup Hakobyan:2008kua. The simulation was tuned to provide a reasonable description of the data. The possible sources of systematic uncertainties studied include acceptance effects, event selection, and particle misidentification. For the azimuthal RMS widths and broadenings, there was a small systematic uncertainty due to the finite bin width used to calculate them. Other sources of systematic uncertainty, such as cross-contamination between bins, beam luminosity, trigger efficiency, Coulomb effects, and time-dependent effects, were found to be negligible. We show the contributions of various types of systematic uncertainties for the correlation functions and the ratio $R$ in Table 4 and for the RMS widths and broadenings in Table 4. Each of the listed uncertainties are “point-to-point”, affecting each bin by a different amount (although there may be some correlation between the bins).
The total relative systematic uncertainty ranged from 15-41% for $C(\Delta\phi,\Delta Y)$, 6-21% for $R$, and 3-6% for $\sigma$. For $b$, the systematic uncertainty was about 4% for almost all bins [^4] The one exception is the highest $\Delta Y$ bin for Fe, where both the statistical and systematic uncertainties are much larger than usual. The reason for this is that the measured $\sigma$ values for C and Fe are very close together, much less than one statistical error bar. Therefore, the uncertainties on their quadrature difference are very large.. Details of these studies on systematic uncertainties can be found in the Supplementary Materials supplementary.
@@
Figure 2: Correlation functions for D, C, Fe, and Pb (from top to bottom, first four rows), with $\Delta\phi$ on the $x$ axis within each panel and different bins of $\Delta Y$ for each column. These are compared to calculations from the BeAGLE model Chang:2022hkt (curves). The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature). Bottom row: same for the ratios between the nuclear and deuterium correlation functions.

### v3

@@ Abstract
We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus ($eA$) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark. The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets. For heavier nuclei, the pion-proton correlation function is more spread-out in azimuth than for lighter ones, and this effect is more pronounced in the $\pi p$ channel than in earlier $\pi\pi$ studies. The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets. These trends are qualitatively reproduced by state‑of‑the‑art $eA$ event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing. At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in $eA$ scattering.
@@ 1 Introduction
The production and propagation of hadrons through nuclei remain central topics in quantum chromodynamics (QCD). Key questions include: “How does the process of hadronization manifest within a nucleus?”, and “How are different hadrons produced in a single scattering event correlated with each other?” Nuclear Science Advisory Committee . Addressing these questions calls for detailed studies of diverse reaction channels, encompassing multiple hadron flavor combinations in the final state and spanning a broad kinematic range, to probe the different timescales of hadronization.
Despite extensive theoretical and phenomenological work on nucleon production in nuclear deep-inelastic scattering (DIS) Adams and others [1995], Degtyarenko et al. [1997], Matsinos and others [1989], Ciofi degli Atti and Kopeliovich [2005], Palli et al. [2009], Larionov and Strikman [2020], Cai et al. [2019], Chang et al. [2022], Robles Gajardo et al. [2022], no dedicated study has correlated the production of low-energy protons with the kinematics of leading hadrons, whose spectra are known to be modified by cold nuclear matter effects as observed by the HERMES Airapetian and others [2001, 2003, 2007, 2010, 2011] and CLAS Moran and others [2022], Chetry and others [2023], El Fassi and others [2012], Mineeva et al. [2025] collaborations. The present work addresses this gap by employing pion-proton correlations over a wide azimuthal and rapidity range, thereby linking studies of current and nuclear fragmentation in DIS off deuterium and heavy nuclei.
In this study, we examine electron–nucleus ($eA$) DIS events that contain a leading $\pi^{+}$ (which retains most of the energy transferred to the struck quark) accompanied by a secondary proton, as illustrated in Fig. 1. Correlating these two hadrons provides new insight into how the nucleus responds to a fast quark or hadron and how the hadronization process is modified within the nucleus. Most of the observed protons come from fragmentation of the nuclear remnant. Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the fast quark or its hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum. This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements Paul and others [2025].
@@
Protons produced through nuclear fragmentation, historically referred to as “grey” protons, were investigated in earlier scattering experiments with both hadronic and leptonic beams Chemakin and others [1999], Adams and others [1995], Degtyarenko et al. [1997]. Measurements of slow protons, together with model calculations, served to estimate the collision centrality, expressed as the mean number of projectile–nucleon interactions. Lepton measurements offer a cleaner test of intra-nuclear hadron-cascade models; in hadron collisions, multiple projectile–nucleon interactions complicate particle‐production mechanisms. Similar studies were carried out with neutrino DIS on nuclear targets Matsinos and others [1989], although with much more limited precision.
Ciofi degli Atti et al. Ciofi degli Atti and Kopeliovich [2005], Palli et al. [2009] proposed that the dependence of slow-proton production over a wide kinematic range can serve to discriminate among hadronization models, and their calculation reproduced well the data from Ref. Adams and others [1995]. Larionov et al. Larionov and Strikman [2020] showed that slow-neutrons provide similar insight into intranuclear cascades.
The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi and others [2016], Abdul Khalek and others [2022], since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in $eA$ collisions Abdul Khalek and others [2022], Chang et al. [2022], Robles Gajardo et al. [2022].
Further, the measurement of $\pi p$ correlations in $eA$ DIS can be used as a reference for analogous neutrino measurements, for instance at MINERvA Aliaga and others [2014], which overlaps the kinematics of the measurements presented in this work, or with legacy data from NOvA Ayres et al. [2007]. It was suggested in Ref. Cai et al. [2019] that left-right asymmetries in $\pi p$ correlations could be measured in $\nu A$ scattering to probe pion-absorption effects.
@@ 2 Experimental setup
The CLAS Mecking and others [2003] detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters (CC), and an electromagnetic calorimeter (EC). Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter. The drift chambers functioned in a toroidal magnetic field up to 2 Tm, where the magnet’s polarity made negatively charged particles bend toward the beam pipe. The resulting resolutions for charged particles were $\sigma_{p}/p=0.5\%$ for the momentum, 2 mrad for the polar angle, and 4 mrad for the azimuthal angle. The data were selected with a trigger that required at least one electron candidate with momentum $p>500$ MeV/c/c.
@@ 3 Event selection and observables
Following Refs. El Fassi and others [2012], Moran and others [2022], Paul and others [2022, 2025], electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC. Background from $\pi^{-}$ was suppressed to the $1\%$ level using the CC and the EC. Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs. The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the CC to suppress proton background. Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution. The protons’ energies were corrected for energy loss in the target using the same procedure as was used in Ref. Chetry and others [2023].
Similar to Refs. Moran and others [2022], Paul and others [2022, 2025], we selected events with $Q^{2}>1$ $2W>2 GeV/c2c^{2}$ and $2.3<\nu<4.2$ GeV. Here, $Q^{2}$ is the square of the four-momentum transfer from the electron to the struck nucleon,
@@
is the magnitude of the four-momentum of the virtual photon+struck nucleon system, where $m_{p}$ is the mass of a proton, $\nu=E-E^{\prime}$ is the energy transfer, and $E$ and $E^{\prime}$ are the beam- and scattered-electron energies. The cuts on $Q^{2}$ and $W$ are conventional for DIS measurements. The minimum cut on $\nu$ is just above the minimum value of $\nu$ possible with these cuts on $Q^{2}$ and $W$. The maximum value of $\nu$ corresponds to a cut on the inelasticity, $y\equiv\nu/E<0.84$.
We selected events with a “leading” pion, defined as having fractional energy $z_{1}=E_{h}/\nu>0.5$ Airapetian and others [2006], Paul and others [2022, 2025], where $E_{h}$ is the energy of the pion. The purpose of this cut is to increase the likelihood that the pion originated from the struck quark and to avoid ambiguity on which pion in the event is considered “leading”. Following Ref. Hen and others [2014], we required the proton to have momentum greater than 350 MeV/c/c, since the proton detection efficiency drops off below this threshold. The maximum proton momentum in our accepted $\pi p$ event sample was about 2.7 GeV/c/c. In order to ensure that the azimuthal angles of both particles can be reconstructed with reasonable resolution, we further required that both the proton and the leading pion have transverse momenta (defined with respect to the momentum-transfer vector $\vec{q}$) greater than 70 MeV/c/c.
@@
where $E_{h}$ is the hadron’s energy and $p_{z,h}$ is its momentum in the momentum-transfer direction. This range in $\Delta Y$ contains nearly 99% of all $\pi^{+}p$ pairs, and allows for a convenient for binning in $\Delta Y$. This differs from the $\Delta Y$ range in the di-pion analysis in Ref. Paul and others [2025], where the observed range in $\Delta Y$ for $\pi^{+}\pi^{-}$ was approximately -0.5 to 2.5.
We selected particles arising from scattering from either the deuterium or nuclear targets by using the longitudinal vertex position defined by intersecting their trajectories with the beamline. The resulting vertex resolution ensured negligible ambiguity in the target tagging Moran and others [2022].
@@
where $N_{e^{\prime}\pi}$ is the number of events with a scattered electron and a leading $\pi^{+}$, and $\Delta\phi\equiv\phi_{\pi}-\phi_{p}$ is the azimuthal separation between the pion and proton, as defined around the momentum-transfer axis. We determined the expression in parentheses by taking $N_{e^{\prime}\pi p}$, the number of events with a scattered electron, a leading $\pi^{+}$ and a proton in each bin in $|\Delta\phi|$ and $\Delta Y$, and dividing that by the bin widths in both variables ($\pi/8$ and 0.5, respectively), and by a factor of two, since each $|\Delta\phi|$ bin covers two $\Delta\phi$ bins. These correlation functions are corrected for pair-acceptance effects using Monte Carlo (MC) simulations. Following Ref. Paul and others [2025], we also used a data-driven method to correct for contamination of the deuterium sample due to scattering off the aluminum endcap walls of the liquid deuterium cells. Details for both of these corrections can be found in the Supplementary Materials 1.
@@
$C_{0}=\frac{N^{D}_{e^{\prime}\pi}}{N^{D,\rm tot}_{e^{\prime}\pi p}},$ (4)
where $N^{D}_{e^{\prime}\pi}$ is the number of events from the deuterium dataset with the scattered electron and leading pion, and $N^{D,\rm tot}_{e^{\prime}\pi p}$ is the total number of events in the deuterium dataset with the scattered electron, leading pion, and proton, for all kinematic bins. The purpose of this normalization is to make the correlation function an observable that is insensitive to the absolute scale of the single-proton efficiency of the detector systems.
From here, we determined the nuclear-to-deuterium ratio of the correlation functions, $R=C_{A}/C_{D}$, which can be used to compare the correlation functions of the nuclear targets with those of deuterium. Since the same $C_{0}$ value is used for nuclear and deuterium targets, it cancels out in the ratio, which is then equivalent to
$R\equiv\frac{N^{A}_{e^{\prime}\pi^{+}p}/N^{A}_{e^{\prime}\pi^{+}}}{N^{D}_{e^{\prime}\pi^{+}p}/N^{D}_{e^{\prime}\pi^{+}}}.$ (5)
We determined from these correlation functions the azimuthal width $\sigma$, defined by Paul and others [2025]
@@
as well as the azimuthal broadening Paul and others [2025] (using carbon as our reference target [^2] We chose carbon instead of deuterium (as was done in Ref. Paul and others [2025]) for the reference target here, since the results for the nuclear targets are more similar to one another than they are to deuterium.),
@@ 4 Uncertainties
We performed studies on various possible sources of systematic uncertainties using data and simulation studies. For the simulation studies we used the PYTHIA6 event generator and the GSIM package Wolin [1996], which is based on GEANT3 Brun et al. [1994] [^3] It should be noted that Geant4 Agostinelli and others [2003] had not been released yet when GSIM was being developed, and that GEANT3 was still the standard at the time that the data were taken., to simulate the response of the CLAS detector and dual-target setup Hakobyan [2008]. The simulation was tuned to provide a reasonable description of the data. The possible sources of systematic uncertainties studied include acceptance effects, event selection, and particle misidentification. For the azimuthal RMS widths and broadenings, there was a small systematic uncertainty due to the finite bin width used to calculate them. Other sources of systematic uncertainty, such as cross-contamination between bins, beam luminosity, trigger efficiency, Coulomb effects, and time-dependent effects, were found to be negligible. We show the contributions of various types of systematic uncertainties for the correlation functions and the ratio $R$ in Table 4 and for the RMS widths and broadenings in Table 4. Each of the listed uncertainties are “point-to-point”, affecting each bin by a different amount (although there may be some correlation between the bins).
The total relative systematic uncertainty ranged from 15-41% for $C(\Delta\phi,\Delta Y)$, 6-21% for $R$, and 3-6% for $\sigma$. For $b$, the systematic uncertainty was about 4% for almost all bins [^4] The one exception is the highest $\Delta Y$ bin for Fe, where both the statistical and systematic uncertainties are much larger than usual. The reason for this is that the measured $\sigma$ values for C and Fe are very close together, much less than one statistical error bar. Therefore, the relative uncertainties on their quadrature difference are very large.. Details of these studies on systematic uncertainties can be found in the Supplementary Materials 1.
@@
Figure 2: Correlation functions for D, C, Fe, and Pb (from top to bottom, first four rows), with $\Delta\phi$ on the $x$ axis within each panel and different bins of $\Delta Y$ for each column. These are compared to calculations from the BeAGLE model Chang et al. [2022] (curves). The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature). Bottom row: same for the ratios between the nuclear and deuterium correlation functions.